ROTATIONAL SPECTROSCOPY AND STRUCTURE OF THE GLYCIDOL-WATER HYDROGEN-BONDED COMPLEX

\underline{MICHAEL J. TUBERGEN}, ANDREW R. CONRAD, NARDOS H. TEUMELSAN, PENG WANG, Department of Chemistry, Kent State University, Kent, OH 44242.

Glycidol, also known as oxiranemethanol, is known to exist in two different conformational structures with different intramolecular hydrogen bonds from the alcohol to the oxirane oxygen. ^{a} ^{b} We have recorded 22 a− and b−type rotational transitions of the glycidol-water complex; the rotational constants of the complex were determined to be A = 3902.331 MHz, B = 2763.176 MHz, and C = 1966.863 MHz by fitting to the Watson A-reduction Hamiltonian. The moment-of-inertia data was used to fit the structure of the glycidol-water complex. Water was found to insert itself into the hydrogen bonding network of the monomer, making a new network: alcohol to water oxygen and water to oxirane oxygen. MP2/6-311G** ab initio calculations were also used to model the structure and determine the relative conformer energies.

aW. V. F. Brooks and K. V. L. N. Sastry, Can. J. Chem. 53, 2247 (1975)

bK. -M. Marstokk, H. Møllendal, Y. Stenstrøm, Acta Chem. Scand. 46, 432 (1992).